# Scaling Population-Based Reinforcement Learning with GPU Accelerated Simulation

### v2

@@ Abstract
In recent years, deep reinforcement learning (RL) has shown its effectiveness in solving complex continuous control tasks like locomotion and dexterous manipulation. However, this comes at the cost of an enormous amount of experience required for training, exacerbated by the sensitivity of learning efficiency and the policy performance to hyperparameter selection, which often requires numerous trials of time-consuming experiments. This work introduces a Population-Based Reinforcement Learning (PBRL) approach that exploits a GPU-accelerated physics simulator to enhance the exploration capabilities of RL by concurrently training multiple policies in parallel. The PBRL framework is applied to three state-of-the-art RL algorithms – PPO, SAC, and DDPG – dynamically adjusting hyperparameters based on the performance of learning agents. The experiments are performed on four challenging tasks in Isaac Gym – Anymal Terrain, Shadow Hand, Humanoid, Franka Nut Pick – by analyzing the effect of population size and mutation mechanisms for hyperparameters. The results show that PBRL agents achieve superior performance, in terms of cumulative reward, compared to non-evolutionary baseline agents. The trained agents are finally deployed in the real world for a Franka Nut Pick task, demonstrating successful sim-to-real transfer. Code and videos of the learned policies are available on our project website.
Index Terms:
Reinforcement Learning, Dexterous Manipulation, Evolutionary Learning, Sim-to-RealI Introduction
@@
Many domains have seen tremendous advancements of reinforcement learning (RL) applications in recent years, ranging from playing challenging games [1, 2] to learning high-dimensional continuous control in robotics [3, 4, 5]. Tasks such as dexterous manipulation [6], legged locomotion [7], and mobile navigation [8] have been learned using deep RL. A primary challenge in training RL policies is the need for large amounts of training data. RL methods rely on effective exploration to discover control policies, which can be particularly challenging when operating in high-dimensional continuous spaces [9]. Moreover, the performance of the learned policy is highly dependent on the tedious tuning of hyperparameters. Hyperparameter tuning, in fact, can be a very time-consuming process, often requiring many manual trials to determine the best values for the specific task and the learning environment. One way to deal with the problem of data inefficiency is to train in simulation before transferring to reality [10, 11, 12]. However, the time required to train the policy in simulation increases significantly with the task complexity. For example, in [11], learning a block re-orientation task with a robot hand took around 14 days and enormous computing resources. In addition, policies trained in simulation often fail to perform on a real system due to discrepancies between the simulation and the real world. Recent advances in GPU-accelerated simulation, such as Isaac Gym [13, 14], have made it possible to run thousands of parallel environments on a single GPU, which reduces the training times significantly. However, successfully training RL policies still requires carefully tuned hyperparameters to explore efficiently.
@@ I-A2 Population-Based Reinforcement Learning
Training various RL agents provides a mechanism for meta-optimization, utilizing the potential of both learning and evolution [25]. One successful example of PBRL methods is population-based training (PBT) [26], which allows training multiple policies concurrently to enhance the exploration capabilities of the agents in generating diverse behaviors. In PBT, the faster inner loop of policy optimization is ultimately evaluated by the slower-running outer loop of evolutionary selection. PBT trains a population of agents with different hyperparameters and uses a genetic algorithm to update the population periodically. Recently, DexPBT, a decentralized PBRL approach has been proposed to learn dynamic manipulation between two hand-arm systems using parallel simulations [27]. The authors developed a decentralized implementation to evolve agents in distributed computing environments using on-policy RL, achieving impressive results in dexterous manipulation. However, sim-to-real transfer has not been performed, highlighting the complexity of deploying policies on real systems.
@@ I-A3 Sim-to-Real Transfer
Despite the calibration efforts to model the physical system accurately, simulation is still a rough approximation. The differences between the dynamics of simulated and real systems cause a “reality gap” that makes it unlikely for a simulation-trained policy to successfully transfer to a physical system.
In literature, researchers have put a significant effort into diminishing this gap: to this aim, most of the approaches leverage domain randomization [4, 6, 16, 17, 28, 29] to expose the policy to a wide range of observation distributions in simulation, thus improving generalization onto a real system. Nevertheless, naive domain randomization might not be sufficient to completely attenuate the dynamics gap: for instance, [30] employs a specific network to mimic the real actuation system. Another technique in this context is policy-level action integrator (PLAI) [10], a simple yet effective algorithm aimed at compensating the sim-to-real dynamic discrepancies with an integral action, which proved to be paramount for a successful transfer.
@@ I-B Contribution
This paper investigates a population-based reinforcement learning (PBRL) framework in robotics that allows the training of a population of agents by exploiting GPU-based massively parallel simulation in order to tune the hyperparameters during training. The PBRL framework is evaluated on four complex tasks in locomotion and manipulation available in Isaac Gym [13]: Anymal Terrain, Humanoid, Shadow Hand, and Franka Nut Pick (Figure 1). The results show that better performance is achieved when training a population of agents compared to a single-agent baseline on all tasks. The comparison is provided across 3 RL algorithms (PPO, SAC, and DDPG), varying the number of agents in a population, and across different hyperparameter mutation schemes. Finally, the PBRL agents are deployed on a real Franka Panda robot for a Franka Nut Pick task, without any policy adaptation phase on the physical system. In summary, the main contributions of this work are:
@@ II-B Population-Based Training
In standard RL, the agent aims to learn an optimal policy by interacting with an environment and iteratively updating the policy through some kind of optimization method. In contrast, PBRL uses a population of agents , each interacting with the environment independently to collect experience and learn its own policy. Using evolutionary selection, the population is evaluated periodically based on a fitness metric, and best-performing members replace the worst-performing members, i.e., weights of the best agents are copied over, along with the mutated hyperparameters.
@@ III-B Results
The experiments focus on optimizing the hyperparameters of the RL agents in a population and comparing the results against non-evolutionary baseline agents. For each case of baseline agents, 4 experiments are run with different seeds. Table I and II provide the hyperparameters for on-policy and off-policy algorithms, listing the sampling ranges of those optimized through the PBRL Algorithm 1. The initial values for each agent are sampled uniformly from a prior distribution with a given range.
@@ III-B1 PBRL-PPO
III-B2 PBRL-SAC
The hyperparameters optimized in PBRL-SAC include the LR of the actor-critic networks and the target entropy factor to control the exploration behavior. Entropy is a central theme in SAC agents as the policy is trained to maximize the trade-off between the expected return and the entropy. The experiments in PBRL-SAC are run with a population size of . The replay buffers in off-policy methods require more memory allocation on GPU than on-policy methods. Since all agents in PBRL-SAC train with their individual replay buffers, the maximum population size is set to 8 due to the limited GPU memory. The training performance of SAC and PBRL-SAC is shown in Figure 4. PBRL-SAC improves the training performance compared to non-evolutionary SAC on 3 out of 4 tasks, yielding a remarkable improvement on both Shadow Hand and Franka Nut Pick, while comparable results are achieved on Humanoid, probably due to the limited task complexity.
III-B3 PBRL-DDPG
@@ III-C Sim-to-Real Transfer
In the real experiments, we replicate the Franka Nut Pick task [4] by deploying a PBRL-PPO policy, without any real-world adaptation on a physical robot, deploying the actions with PLAI [10]. The robot detects the nuts utilizing Mask-RCNN [32], fine-tuned on real-world images captured with a wrist-mounted Intel RealSense D435 camera, using the IndustRealLib codebase [10].
Compared to the original task [4], we applied the following changes to make the simulated environment as close as possible to the real setup: (i) employing a Task-Space Impedance (TSI) controller [33] instead of an Operational-Space motion Controller (OSC) [34] to comply with the actual low-level controller111The control laws are specified in [4] and in reference works [33, 34]; (ii) randomizing the nut’s initial position to resemble the actual robot workspace; (iii) changing the observation space to include the 7-DOF joint configuration, the measured end-effector pose, and the estimated nut pose. The parameters used in the simulated environment and the real controller are reported in Table IV.
During experiments, the following policies were deployed, performing 30 real-world trials of Franka Nut Pick task for each policy: (i) 2 agents from a population of 8, trained with PBRL-PPO, specifically the “best” and the “worst” agent; (ii) the “best” agent trained with baseline PPO. With “best” and “worst” we indicate the agents achieving the highest and lowest success rate in simulation, where success is defined as reaching, grasping, and lifting the nut, without losing contact during the lifting phase. As shown in Figure 7, PBRL-PPO with has the highest success rate. Remarkably, even the success rate of the worst agent in is higher than that of the best PPO agent ( vs. ), also highlighting lower variance compared to both baseline PPO and PBRL-PPO with .
Deploying both PPO and PBRL-PPO agents onto a real robot leads to task completion (shown in Figure 8), yet with different success rates, as summarized in Table V. Particularly, both PBRL-PPO agents yield higher success rates than PPO, with the “best” agent performing better than the “worst” one, indeed confirming the ranking attained in simulation. This demonstrates that PBRL-PPO, favoring exploration across the agents, yields a successful real-world transfer, despite reaching similar rewards as the standard PPO. Informally, the best PBRL-PPO agent also exhibited recovery behavior during task execution after perturbation by the human.
@@ III-D Discussion
While the PBRL agents perform better than the non-evolutionary agents in almost all the experiments, the impact of population size across RL algorithms and tasks shows no consistent pattern. One may hypothesize that larger and more diverse populations might lead to a better final performance. However, the results in this work indicate that using a larger population size does not necessarily yield substantial benefits for every task. This is in contrast to the common belief that population-based methods rely on larger population sizes to effectively explore the hyperparameter space [26, 35]. The optimal population size, instead, depends on various factors, including task complexity, RL algorithm, and interaction dynamics among agents. While larger populations offer increased exploration potential, they also suffer from diminished exploitation capabilities due to increased competition, leading to lower performance in less challenging tasks where smaller populations suffice. Perhaps, larger population sizes also perform better when the task complexity gradually increases requiring extensive exploration as in Anymal Terrain, which implements curriculum learning.
@@ IV-A Summary
In this paper, a PBRL framework has been employed for training a population of RL agents by making use of high-throughput GPU-accelerated simulation. The first simulation results of PBRL using on-policy and off-policy methods are provided on a series of benchmark tasks proposed in [13] by investigating the effect of population size and different mutation schemes. The results showed the effectiveness of PBRL in improving final performance by adapting the hyperparameters online. PBRL agents have been deployed on real hardware for the first time, demonstrating smooth and successful transfer, without any policy adaptation or fine-tuning. Finally, we release the codebase to train PBRL agents and hope that it will empower researchers to further explore and extend the capabilities of PBRL algorithms on challenging robotic tasks.
@@ IV-B Future Works
Many interesting directions remain for future research. An immediate extension could be to use a dedicated fitness metric for each population sub-group to prioritize long-term performance [36]. This can circumvent a greedy decision process of population-based methods that may lead to undesirable behavior later in the training.
Additionally, it remains to be seen how the PBRL agents perform on contact-rich tasks (e.g., dexterous manipulation, and assembly) in the real world with sim-to-real techniques. With a number of agents training in parallel environments, the PBRL framework has the potential to solve complex robotic manipulation tasks, making them feasible and computationally tractable, accounting for diverse skills labeling and learning.

### v3

@@ Abstract
In recent years, deep reinforcement learning (RL) has shown its effectiveness in solving complex continuous control tasks like locomotion and dexterous manipulation. However, this comes at the cost of an enormous amount of experience required for training, exacerbated by the sensitivity of learning efficiency and the policy performance to hyperparameter selection, which often requires numerous trials of time-consuming experiments. This work introduces a Population-Based Reinforcement Learning (PBRL) approach that exploits a GPU-accelerated physics simulator to enhance the exploration capabilities of RL by concurrently training multiple policies in parallel. The PBRL framework is applied to three state-of-the-art RL algorithms – PPO, SAC, and DDPG – dynamically adjusting hyperparameters based on the performance of learning agents. The experiments are performed on four challenging tasks in Isaac Gym – Anymal Terrain, Shadow Hand, Humanoid, Franka Nut Pick – by analyzing the effect of population size and mutation mechanisms for hyperparameters. The results demonstrate that PBRL agents outperform non-evolutionary baseline agents across tasks essential for humanoid robots, such as bipedal locomotion, manipulation, and grasping in unstructured environments. The trained agents are finally deployed in the real world for the Franka Nut Pick manipulation task. To our knowledge, this is the first sim-to-real attempt for successfully deploying PBRL agents on real hardware. Code and videos of the learned policies are available on our project website.
I Introduction
@@
Many domains have seen tremendous advancements of reinforcement learning (RL) applications in recent years, ranging from playing challenging games [1, 2] to learning high-dimensional continuous control in robotics [3, 4, 5]. Tasks such as dexterous manipulation [6], legged locomotion [7], and mobile navigation [8] have been learned using deep RL. A primary challenge in training RL policies is the need for large amounts of training data. RL methods rely on effective exploration to discover control policies, which can be particularly challenging when operating in high-dimensional continuous spaces [9]. Moreover, the performance of the learned policy is highly dependent on the tedious tuning of hyperparameters. Hyperparameter tuning can be a very time-consuming process, often requiring many manual trials to determine the best values for the specific task and the learning environment. One way to deal with the problem of data inefficiency is to train in simulation before transferring to reality [10, 11, 12]. However, the time required to train the policy in simulation increases significantly with the task complexity. For example, in [11], learning a block re-orientation task with a robot hand took around 14 days and enormous computing resources. In addition, policies trained in simulation often fail to perform on a real system due to discrepancies between the simulation and the real world. Recent advances in GPU-accelerated simulation, such as Isaac Gym [13, 14], have made it possible to run thousands of parallel environments on a single GPU, which reduces the training times significantly. However, successfully training RL policies still requires carefully tuned hyperparameters to explore efficiently.
@@ I-A2 Population-Based Reinforcement Learning
Training various RL agents provides a mechanism for meta-optimization, utilizing the potential of both learning and evolution [25]. One successful example of PBRL methods is population-based training (PBT) [26], which allows training multiple policies concurrently to enhance the exploration capabilities of the agents in generating diverse behaviors. PBT trains a population of agents with different hyperparameters and uses a genetic algorithm to update the population periodically. Recently, DexPBT, a decentralized PBRL approach has been proposed to learn dynamic manipulation between two hand-arm systems using parallel simulations [27]. The authors developed a decentralized implementation to evolve agents in distributed computing environments using on-policy RL, achieving impressive results in dexterous manipulation. However, sim-to-real transfer has not been performed, highlighting the complexity of deploying policies on real systems.
@@ I-A3 Sim-to-Real Transfer
Despite the calibration efforts to model the physical system accurately, simulation is still a rough approximation. The differences between the dynamics of simulated and real systems cause a “reality gap” that makes it unlikely for a simulation-trained policy to successfully transfer to a physical system. In literature, researchers have put a significant effort into diminishing this gap: to this aim, most of the approaches leverage domain randomization [4, 6, 16, 17, 28, 29] to expose the policy to a wide range of observation distributions in simulation, thus improving generalization onto a real system. Nevertheless, naive domain randomization might not be sufficient to completely attenuate the dynamics gap: for instance, [30] employs a specific network to mimic the real actuation system. Another technique in this context is policy-level action integrator (PLAI) [10], a simple yet effective algorithm aimed at compensating the sim-to-real dynamic discrepancies with an integral action, which proved to be paramount for a successful transfer.
@@ I-B Contribution
This paper investigates a population-based reinforcement learning (PBRL) framework in robotics that allows the training of a population of agents by exploiting GPU-based massively parallel simulation to dynamically adjust the hyperparameters during training. We evaluate the PBRL framework on four complex tasks that require learning essential skills for humanoid robots: Anymal Terrain, Humanoid, Shadow Hand, and Franka Nut Pick (Figure 1), available in Isaac Gym [13]. The results show that better performance is achieved when training a population of agents compared to a single-agent baseline on all tasks. The comparison is provided across 3 RL algorithms (PPO, SAC, and DDPG), varying the number of agents in a population, and across different hyperparameter mutation schemes. Finally, the PBRL agents are deployed on a real Franka Panda robot for a Franka Nut Pick task, without any policy adaptation phase on the physical system. In summary, the main contributions of this work are:
@@ II-B Population-Based Training
In standard RL, the agent aims to learn an optimal policy by interacting with an environment and iteratively updating the policy through some kind of optimization method. In contrast, PBRL uses a population of agents , each interacting with the environment independently to collect experience and learn its own policy. Using evolutionary selection, the population is periodically evaluated based on a fitness metric, and best-performing members replace the worst-performing members, i.e., weights of the best agents are copied over, along with the mutated hyperparameters.
@@ III-B Results
The experiments focus on optimizing the hyperparameters of the RL agents in a population and comparing the results against non-evolutionary baseline agents. For each case of baseline agents, 4 experiments are run with different seeds. Tables I and II list the hyperparameters for on-policy and off-policy algorithms, including the sampling ranges of those optimized through the PBRL Algorithm 1. The initial values for each agent are uniformly sampled from a prior distribution with a given range.
@@ III-B1 PBRL-PPO
III-B2 PBRL-SAC
In PBRL-SAC, the optimized hyperparameters include the LR of the actor-critic networks and the target entropy factor. Entropy is key in SAC agents as the policy is trained to maximize the trade-off between the expected return and exploration. Experiments are run with a population size of . Due to higher memory needs for replay buffers in off-policy methods, the maximum population size is limited to 8. The training performance of SAC and PBRL-SAC is shown in Figure 4. PBRL-SAC improves the training performance compared to non-evolutionary SAC on 3 out of 4 tasks, yielding a remarkable improvement on both Shadow Hand and Franka Nut Pick, while comparable results are achieved on Humanoid, probably due to the limited task complexity.
III-B3 PBRL-DDPG
@@ III-C Sim-to-Real Transfer
In the real experiments, we replicate the Franka Nut Pick task [4] by deploying a PBRL-PPO policy, without any real-world adaptation, executing the actions with PLAI [10]. The robot detects the nuts utilizing Mask-RCNN [32], fine-tuned on real-world images captured with a wrist-mounted Intel RealSense D435 camera, using the IndustRealLib codebase [10]. Compared to the original task [4], we applied the following changes to make the simulated environment resemble real setup: (i) employing a Task-Space Impedance (TSI) controller [33] instead of an Operational-Space motion Controller (OSC) [34] to comply with the actual low-level controller111The control laws are specified in [4] and in reference works [33, 34]; (ii) randomizing the nut’s initial position to reflect the actual robot workspace; (iii) changing the observation space to include the 7-DOF joint configuration, the measured end-effector pose, and the estimated nut pose. The parameters used in the simulated environment and the real controller are reported in Table IV.
During experiments, the following policies were deployed, performing 30 real-world trials of Franka Nut Pick task for each policy: (i) 2 agents from a population of 8, trained with PBRL-PPO, specifically the “best” and the “worst” agent; (ii) the “best” agent trained with baseline PPO. With “best” and “worst” we indicate the agents achieving the highest and lowest success rate in simulation, where success is defined as reaching, grasping, and lifting the nut, without losing contact during the lifting phase. As shown in Figure 7, PBRL-PPO with has the highest success rate. Remarkably, even the success rate of the worst agent in is higher than that of the best PPO agent ( vs. ).
Deploying both PPO and PBRL-PPO agents onto a real robot leads to task completion (shown in Figure 8), yet with different success rates, as summarized in Table V. Particularly, both PBRL-PPO agents yield higher success rates than PPO, with the “best” agent performing better than the “worst” one, indeed confirming the ranking attained in simulation. Unlike the baseline PPO agent, which continued to produce small movements after reaching the target, PBRL-PPO agents remained more stable, leading to a higher success rate. This demonstrates that PBRL agents, while achieving similar rewards to a single agent, learn behaviors that exhibit greater robustness to environment variability due to the diversity in agent populations. Informally, the best PBRL-PPO agent also exhibited recovery behavior during task execution after perturbation by the human.
@@ III-D Discussion
While the PBRL agents perform better than the non-evolutionary agents in almost all the experiments, the impact of population size across RL algorithms and tasks shows no consistent pattern. One may hypothesize that larger and more diverse populations might lead to a better final performance. However, the results in this work indicate that using a larger population size does not necessarily yield substantial benefits for every task. This is in contrast to the common belief that population-based methods rely on larger population sizes to effectively explore the hyperparameter space [26, 35]. The optimal population size, instead, depends on various factors, including task complexity, RL algorithm, and interaction dynamics among agents. While larger populations offer increased exploration potential, they also suffer from diminished exploitation capabilities due to increased competition, leading to lower performance in less challenging tasks where smaller populations suffice. Larger population sizes seem to perform better when the task complexity gradually increases requiring extensive exploration as in Anymal Terrain, which implements curriculum learning.
@@ IV-A Summary
In this paper, a PBRL framework has been employed to train a population of RL agents by making use of high-throughput GPU-accelerated simulation. The first simulation results of PBRL using on-policy and off-policy methods are provided on a series of locomotion and manipulation benchmark tasks proposed in [13] by investigating the effect of population size and different mutation schemes. The results showed the effectiveness of PBRL in improving final performance through online adaptation of hyperparameters. PBRL agents have been deployed on real hardware for the first time, demonstrating smooth and successful transfer, without any policy adaptation or fine-tuning. Finally, we released the codebase to train PBRL agents and hope that it will empower researchers to further explore and extend the capabilities of PBRL algorithms.
@@ IV-B Future Works
Many interesting directions remain for future research. An immediate extension could be to use a dedicated fitness metric for each population sub-group to prioritize long-term performance [36]. This can circumvent a greedy decision process of population-based methods that may lead to undesirable behavior later in the training. Additionally, it remains to be seen how the PBRL agents perform on contact-rich tasks (e.g., dexterous manipulation or assembly) in the real world with sim-to-real techniques. With a number of agents training in parallel environments, the PBRL framework has the potential to solve complex robotic manipulation tasks, making them feasible and computationally tractable.